Earley Corporation issued perpetual preferred stock with a 9% annual dividend. The stock currently yields 6%, and its par value is $100. Round your answers to the nearest cent.

  1. What is the stock's value?
    $  
  2. Suppose interest rates rise and pull the preferred stock's yield up to 11%. What is its new market value?
    $  

9.11
Answer #1

Annual dividend=100*9%=$9

Stock value=Annual dividend/Current yield

a.Stock value=9/0.06=$150

b.Stock value=9/0.11=$81.82(Approx).